solve by using teh multiplication principle to clear fractions
1/5a-1/5= -6
The question is not in the book, it is a review for a test.

Answer by themathprof(21)

solve by using teh multiplication principle to clear fractions
1/5a-1/5= -6
the 2 denominators are 5a, and 5
the LCM is 5a
multiply all THREE terms by 5a to clear fractions
[5a]1/5a-1[5a]/5= -6[5a]
1-1a=-30a
1=-30a+1a
1=-29a
-1/29=a
See problems like this and lots of
word problems worked out step by
step in my FREE math911 tutorials
software.
Download from www.math911.com
Tell your friends, classmates & teacher
Professor Martin Weissman
mathprof@hotmail.com